Ratio and Proportion
1. A lawyer bills her clients $200 per hour of service. If a client's case requires 39 hours to complete, use proportion to calculate how much the client will owe the lawyer (excluding tax).
I am having a really hard time figuring out this word problem. I have always had trouble figuring them out because I never know where to begin. I figured it I multiply 39*200=7,800.00, but I know this is totally wrong. Can you please help me! I have two more word problems to do that are similar to this. Thank you so much for assisting.

Answer by stanbon(75887)

A lawyer bills her clients $200 per hour of service. If a client's case requires 39 hours to complete, use proportion to calculate how much the client will owe the lawyer (excluding tax).
------------------------
hrs/charge = hrs/charge
--------------------------
1/200 = 39/x
Cross-multiply to get:
x = 39*200
x = $7800
---
You are totally right.
===============
Cheers,
Stan H.